find dy/dx when
y=x^x + x^1/x

Grade:12

1 Answers

Revathi Mukkamala
9 years ago
Let u = x^x
log u = x log x
differentiating wrt x we get du/dx
Let v = x^1/x
log v = (1/x) log x
diff we get. dv/dx
y = u + v
dy/dx = du/dx + dv/dx